A 2.0kg ball and a 4.0kg ball are connected by a 3.0m long rigid, massless rod. The rod and balls are rotating clockwise about its center of gravity at 16rpm .

What torque will bring the balls to a halt in 9.0s ? Express your answer using two significant figures. Assume the counterclockwise rotation as positive.
Answer #1

Here ,

m1 = 2 Kg

m2 = 4 kg

Here ,

Xcom = 2*3/(2+4)

Xcom = 1 m

Here ,

I = 2*2^2 + 4*1^2

I = 12 Kg.m/s

the moment of inertia is 12 Kg.m/s

Here , wf = 16 rpm

wf = 1.675 rad/s

Now , Using second law of motion

Torque , T = 12 * (-1.675 + 0)/9

T = -2.23 N.m

the torque required is -2.23 N.m